What is the heat change (KJ) when 2.31 moles of chlorine reacts with methane?


1
Expert's answer
2022-04-14T15:19:03-0400

CH4 + Cl2 = CH3Cl + HCl, "\Delta" H = -104 kJ (by the chemical literature)

1 mole correspond to (-104) kJ of the heat change, then 2.31 mole correspond x kJ.

x = 2.31 * (-104)/1 = -240.2 kJ.

Answer: -240.2 kJ.